# Language-Driven Closed-Loop Grasping with Model-Predictive Trajectory Replanning

### v1

@@ Abstract
Combining a vision module inside a closed-loop control system for a seamless movement of a robot in a manipulation task is challenging due to the inconsistent update rates between utilized modules. This task is even more difficult in a dynamic environment, e.g., objects are moving. This paper presents a modular zero-shot framework for language-driven manipulation of (dynamic) objects through a closed-loop control system with real-time trajectory replanning and an online 6D object pose localization. We segment an object within by leveraging a vision language model via language commands. Then, guided by natural language commands, a closed-loop system, including a unified pose estimation and tracking and online trajectory planning, is utilized to continuously track this object and compute the optimal trajectory in real-time. Our proposed zero-shot framework provides a smooth trajectory that avoids jerky movements and ensures the robot can grasp a non-stationary object. Experiment results exhibit the real-time capability of the proposed zero-shot modular framework for the trajectory optimization module to accurately and efficiently grasp moving objects, i.e., up to update rates for the online 6D pose localization module and update rates for the receding-horizon trajectory optimization. These advantages highlight the modular framework’s potential applications in robotics and human-robot interaction; see the video in https://www.acin.tuwien.ac.at/en/6e64/.
@@ I INTRODUCTION
Integrating artificial intelligence (AI) in robotic systems transforms human-robot collaboration, particularly in dynamic and complex environments. Traditional robotic systems typically rely on pre-programmed tasks and controlled environments with limited variability. However, real-world applications demand higher adaptability and precision, especially when robots need to interact with humans in real-time. Imagine we want an assistant robot to grasp a hammer among a clutter of workshop objects such as scissors, scales, screwdrivers, or handrails. The robot must be capable of continuously detecting and grasping the hammer, even if the hammer can be moved. While humans can naturally understand how to grasp a hammer when instructed, it is still challenging for robots to identify precise grasping actions for objects based on natural language commands, a task known as language-driven grasp detection [1]. This simple example encompasses three challenges, i.e., natural language understanding, real-time object localization, and real-time trajectory planning. Despite recent advancements, bridging the gap between language, vision, and control in real-world robotic experiments remains a challenging task [2]. Furthermore, recent works, e.g., LLaRP [3], Open X-embodiment [4], often focus on an end-to-end approach relying on pre-defined large skill datasets. This reliance on individual skill acquisition is usually seen as a significant bottleneck for application in various systems, primarily due to mismatches in dynamic models, which result in the jerky motion of the robot. Furthermore, these end-to-end approaches often neglect kinematic and dynamic constraints, leading to unsafe and inaccurate robot motions. These requirements are significant concerns for meeting industrial standards, where precise and reliable operations are critical. Thus, in this work, our focus is creating a zero-shot modular framework utilizing a vast knowledge of foundation models for vision-language understanding and smoothly controlling the robot at a granular action level, bypassing the need for extensive data collection or manual programming of individual basic movements. Also, with the modular fashion, we can integrate these individual system constraints into the trajectory optimization process to ensure that the robot’s movements are safe and precise.
@@
To address the challenge above in language-driven grasping tasks, we propose a modular zero-shot framework comprising (i) a vision-language model to analyze the human prompts and to ground the mentioned object, (ii) a vision module to localize the object in real-time, and (iii) the receding-horizon trajectory optimization to execute a smooth trajectory toward the object. Specifically, we first utilize the open-vocabulary detectors such as NanoOWL [5] and OWL-ViT2 [6] leveraging contrastively pre-train image and text encoders on large-scale image-text data RegionCLIP [7]. This allows us to extract spatial-geometric information of relevant objects from natural language commands. Next, we employ the unified foundation model for 6D object pose localization, FoundationPose [8], to efficiently estimate and continuously track the object’s 6D pose. Given that the object is not stationary, we employ our recent work, the model-predictive trajectory optimization (MP-TrajOpt)[9], to dynamically guide the robot toward the object. An overview of the proposed framework is illustrated in Figure 1, which is presented in detail in Section III.
@@
The remainder of the paper is as follows. After discussing related work in the next section, Section III presents the proposed framework, including three modular modules. Subsequently, experiments and analysis, including comparisons with other approaches and limitations, are discussed in Section IV. Finally, Section V concludes this work and gives some remarks.
@@ II Related Work
Foundation Models for Robotic Manipulation. Various methods have been proposed to integrate foundation models into robotic manipulation [3, 10]. For instance, Ha et al.[11] use large language models (LLMs) as knowledge bases to decide tasks and utilize sampling-based robot planners to create comprehensive manipulation skill datasets. These datasets can then be used to train diffusion policies for diverse manipulation tasks using natural language commands. While the trend of applying foundation models to robotics is growing [12], it still faces significant uncertainties [13]. One major issue is the need for improved functional competence in these models [13], which limits their ability to tackle novel planning problems that require an understanding of real-world dynamics [14]. Additionally, enabling language models to perceive physical environments often involves providing textual descriptions and incorporating vision during decoding [15]. To close the perception-action loop, an embodied language model must also be capable of action, typically achieved through a library of predefined primitives. However, these primitives vary across robotic platforms, often requiring a large dataset to generalize effectively. Unlike previous methods, we leverage pre-trained language models for their extensive open-world knowledge and focus on the complex task of planning for dynamic object manipulation. Our approach incorporates system dynamic constraints, enabling a seamless integration of perception and action. This method addresses the limitations of previous models by enhancing functional competence and facilitating real-time interaction and adaptation.
@@
Language-driven object detection. Traditionally, language-driven object detection methods were designed for limited categories, a.k.a. closed-vocabulary object detection, like DETR [16]. Recently, using large vision-language models (VLMs) like CLIP [17] has opened doors to open-vocabulary object detection. Unlike traditional methods, open-vocabulary object detection methods OWLv2 [6] and RegionCLIP [7] are not restricted to predefined categories. This allows them to excel in tasks like zero-shot object detection and segmentation based on natural language descriptions. This shift towards open-vocabulary object detection [6] extends beyond just identifying objects. It has applications in related areas like affordance detection [18], where the focus is identifying potential actions that can be performed on objects. Inspired by OWLv2, we utilize this method in our framework to enable zero-shot object detection via natural language commands.
@@
Object pose localization. Recent advancements in object pose estimation have focused on developing techniques that handle novel objects without requiring retraining, addressing a key challenge in applying deep learning models to industrial settings. Nguyen et al. [19] introduced GigaPose, simplifying the pose estimation process to a single correspondence match between a query image and a set of templates. More recently, Wen et al. [8] presented FoundationPose, a unified 6D object pose estimation and tracking model, utilizing large-scale synthetic training data and a novel transformer-based architecture. FoundationPose supports both model-based and model-free setups with novel objects. However, manual segmentation of objects in the first frame from pre-recorded videos is still required. Therefore, in this work, we integrate a language-driven object detection module that could enable end-to-end applications from human commands to robotic grasping actions in a closed-loop system.
@@
Real-time trajectory optimization. In recent years, significant interest has been in extending trajectory optimization to online planning using a receding horizon. This interest spans both gradient-based methods [20] and sampling-based approaches [21]. The sequence-of-constraints Model Predictive Control (MPC) introduced by Toussaint et al. [22] addresses task-and-motion planning (TAMP) in three phases: initially, task planning generates waypoints; subsequently, the timing of these waypoints is optimized to form a reference trajectory; finally, MPC tracks this reference trajectory to compute collision-free paths over a short planning horizon. Note that this approach requires a global reference to incorporate waypoints into the MPC. Tackling this issue, in our recent work [9], we propose the model-predictive trajectory optimization (MP-TrajOpt) that avoids calculating a reference trajectory for waypoint timing. Instead, this MPC approach employs a cost-to-go analysis for waypoints. The advantages of integrating MP-TrajOpt are twofold. First, to complete a general grasping task, the robot must navigate specific waypoints to avoid collisions between its gripper and the object. For example, in a top-down grasp, the robot should pass through a waypoint located above the object along the -axis, as illustrated in Figure 3. Second, passing waypoints with non-zero velocity allows for a more natural and smooth grasping action, which differs from most state-of-the-art methods [23, 24].
@@
Leveraging vast knowledge from foundation models, we tackle the language-driven task in this world. Our contributions are summarized as follows:
@@ III Methods
In this section, we explain the three main modules: language-driven object detection, real-time object pose localization, and online trajectory optimization (MP-TrajOpt), illustrated in Figure 1. Consider a scenario in a workshop where a human instructs a robot to grasp a tool with a known CAD model and place it in a predefined location. First, the language-driven object detection module processes the image to determine the 2D location of the tool and generate a binary mask. This mask, combined with the CAD model, is used by the real-time object pose localization module to estimate the initial pose of the tool. The pose is then refined using a linear Kalman filter for smoothness. Finally, the online trajectory optimization module plans and executes the grasp and placement actions, ensuring smooth and collision-free movements.
@@ III-A Language-Driven Object Detection
To determine the 2D location of a target object specified in a natural language command, such as ”Grasp the orange drill,” we first extract key phrases from the command, in this case, ”orange drill.” Concurrently, an RGB image is input into the OWLv2 model [6] to generate pseudo-box annotations for image-text pairs, as illustrated in Figure 2. Although we utilize the pre-trained weights of the OWLv2 model without additional training, this approach is both efficient and effective at pinpointing the object’s center within the image. In comparison to more complex segmentation methods, such as SAM [25] or CNOS [26], which require several seconds for inference, the OWLv2 model can ground the object within across all experimental cases. Upon obtaining the bounding box of the target object, a binary mask is generated. This binary mask serves as the input for the subsequent object pose localization process.
@@ III-B Object pose localization module
In the pose initialization stage, we align and synchronize data from the depth sensor and the RGB sensor of an RGB-D camera. Subsequently, leveraging the object detection mask generated by the method mentioned earlier, we formulate an initial coarse pose estimation. This initial estimation serves as input to a pose refinement network in the subsequent stage. The architecture of this refinement network, depicted in the second stage of Figure 2, is designed to infer translation updates and rotation updates of the corresponding object. Each update is individually processed through a transformer encoder[27] and then linearly projected to the output dimension. Notably, both and are expressed relative to the camera frame. The initial coarse pose is refined utilizing the following update equations [8]:
where indicates the update operation on . This disentangled representation approach effectively decouples the translation update from the updated orientation. The refinement network is trained under supervision [8] using the L2 loss function:
@@
To enhance the accuracy and smoothness of pose estimation ] of the object, we incorporate a simple Kalman Filter (KF) [28] of the state vector
with is the vector of three Euler angles computed from the object’s orientation [29]. We consider the process model for the LKF with the update time and at the time index in the following
@@ III-C1 Mathematical model
with the generalized coordinates , the positive definite mass matrix , the matrix of Coriolis and centrifugal terms , the gravitational forces , and the generalized torques . We consider only a rigid-body model for simplicity of presentation, neglecting elastic joints. For the full elastic joint model, the reader is referred to [30] for a detailed derivation of the model and the controller.
@@ III-C3 Planning
the cost terms and drive the trajectory to the waypoint and the goal, respectively. The regularization cost on ensures smooth trajectories, and is a collision avoidance term. The weights are tuning parameters. The trajectory must adhere to the system dynamics (15b) with the initial states obtained by the previous planner iteration (15c). The endpoint is constraint to an equilibrium point in (15d), and the states and inputs are bounded using upper , and lower , bounds. The endpoints of the split horizon are constrained to a waypoint tolerance set and a goal tolerance set , respectively. For details on computing the horizon lengths and , refer to [9].
We use a state machine to guide the robot through the pick-and-place task. First, the robot must approach the object to grasp it. We derive a grasp pose and a pre-grasp waypoint from the object pose. The pre-grasp pose is placed above the object to ensure a proper approach direction and a collision-free grasp. Figure 3 illustrates a simple example of this pre-grasp waypoint. Note that moving the object further than a tolerance dynamically changes these poses. When the robot end-effector reaches the grasp pose, the robot grasps the object and can plan toward the place location. Similarly to the grasp poses, a pre-place and place pose are defined to place the object.
@@ IV Experiments
Experiments are carried out on grasping tasks in real-world scenarios with a 7-DoF KUKA LBR iiwa 14 R820 and novel objects, captured in Figure 4. We utilize the D435i RealSense camera in an eye-to-hand setup where this camera observes the robot within its workspace like in Figure 3.
Three modules of the proposed framework, see Figure 1, are processed on a desktop with Intel Core i7 10700K and 32 GB of RAM. Inference processes of the language-driven module and the object pose localization are run on the NVIDIA RTX 3080 with 8GB of VRAM. The optimization problem described in Equation (15) is implemented as a Python-based ROS node utilizing CasADi [33]. It is solved using the nonlinear interior point solver IPOPT [34] and the MA57 linear solver, all running on an Ubuntu 22.04 system. This implementation achieves planning times under , which includes the online resolution of the analytic inverse kinematics for new Cartesian waypoints and a desired goal [31]. To ensure compatibility of inverse kinematics solutions for these waypoints, the solution nearest to the previous one is selected using a least-squares criterion.

### v2

@@ Abstract
Combining a vision module inside a closed-loop control system for a seamless movement of a robot in a manipulation task is challenging due to the inconsistent update rates between utilized modules. This task is even more difficult in a dynamic environment, e.g., objects are moving. This paper presents a modular zero-shot framework for language-driven manipulation of (dynamic) objects through a closed-loop control system with real-time trajectory replanning and an online 6D object pose localization. We segment an object within by leveraging a vision language model via language commands. Then, guided by natural language commands, a closed-loop system, including a unified pose estimation and tracking and online trajectory planning, is utilized to continuously track this object and compute the optimal trajectory in real-time. Our proposed zero-shot framework provides a smooth trajectory that avoids jerky movements and ensures the robot can grasp a non-stationary object. Experimental results demonstrate the real-time capability of the proposed zero-shot modular framework to accurately and efficiently grasp moving objects. The framework achieves update rates of up to for the online 6D pose localization module and for the receding-horizon trajectory optimization. These advantages highlight the modular framework’s potential applications in robotics and human-robot interaction; see the video on https://www.acin.tuwien.ac.at/en/6e64/.
@@ I INTRODUCTION
Integrating artificial intelligence (AI) in robotic systems transforms human-robot collaboration, particularly in dynamic and complex environments. Traditional robotic systems typically rely on pre-programmed tasks and controlled environments with limited variability. However, real-world applications demand higher adaptability and precision, especially when robots need to interact with humans in real-time. Imagine we want an assistant robot to grasp a hammer among a clutter of workshop objects such as scissors, scales, screwdrivers, or handrails. The robot must be capable of continuously detecting and grasping the hammer, even if the hammer can be moved. This simple example encompasses three challenges, i.e., natural language understanding, real-time object localization and tracking, and real-time trajectory planning. Despite recent advancements, bridging the gap between language, vision, and control in real-world robotic experiments remains a challenging task [1]. Furthermore, recent works, e.g., LLaRP [2], and Open X-embodiment [3], often focus on an end-to-end approach relying on pre-defined large skill datasets. This reliance on individual skill acquisition is a significant bottleneck for application in various systems, primarily due to mismatches in dynamic models, which result in jerky motions of the robot. Furthermore, these end-to-end approaches often neglect kinematic and dynamic constraints, leading to unsafe and inaccurate robot motions. These requirements are significant concerns for meeting industrial standards, where precise and reliable operations are critical. Thus, in this work, our focus is creating a zero-shot modular framework utilizing a vast knowledge of foundation models for vision-language understanding and smoothly controlling the robot at a granular action level, bypassing the need for extensive data collection or manual programming of individual basic movements. Also, with the modular fashion, we can integrate these individual system constraints into the trajectory optimization process to ensure that the robot’s movements are safe and precise.
@@
To address the challenge above in language-driven grasping tasks, we propose a modular zero-shot framework comprising (i) a vision-language model to analyze the human prompts and to ground the mentioned object, (ii) a vision module to localize and track the object in real-time, and (iii) the receding-horizon trajectory optimization to plan a smooth trajectory toward the object. Specifically, we first utilize open-vocabulary detectors such as NanoOWL [4] and OWLv2 [5] leveraging contrastively pre-trained image and text encoders on the large-scale image-text data set RegionCLIP [6]. The open-vocabulary detector allows us to extract spatial-geometric information of relevant objects from natural language commands. Next, we employ the unified foundation model for 6D object pose localization, FoundationPose [7], to efficiently estimate and continuously track the object’s 6D pose. Given that the object is not stationary, we employ our recent work, the model-predictive trajectory optimization (MP-TrajOpt)[8], to dynamically guide the robot toward the object. Figure 1 illustrates an overview of the proposed framework, described in detail in Section III.
@@
The remainder of the paper is structured as follows. After discussing related work in the next section, Section III presents the proposed framework, including the three modules: language-driven object detection, object pose localization and model-predictive trajectory optimization. Experiments and analysis, including comparisons with other approaches and limitations, are discussed in Section IV. Finally, Section V concludes this work and gives some remarks.
@@ II Related Work
Foundation Models for Robotic Manipulation. Various methods have been proposed to integrate foundation models into robotic manipulation [2, 9]. For instance, Ha et al.[10] use large language models (LLMs) as knowledge bases to decide tasks and utilize sampling-based motion planners to create comprehensive manipulation skill datasets. These datasets can then be used to train diffusion policies for diverse manipulation tasks using natural language commands. While the trend of applying foundation models to robotics is growing [11], it still faces significant uncertainties [12]. One major issue is the need for improved functional competence in these models [12], which limits their ability to tackle novel planning problems that require an understanding of real-world dynamics [13]. Additionally, enabling language models to perceive physical environments often involves providing textual descriptions and incorporating vision during decoding [14]. To close the perception-action loop, an embodied language model must also be capable of action, typically achieved through a library of predefined primitives. However, these primitives vary across robotic platforms, often requiring a large dataset to generalize effectively. Unlike previous methods, we leverage pre-trained language models for their extensive open-world knowledge and focus on the complex task of planning for dynamic object manipulation. Our approach incorporates system dynamic constraints, enabling a seamless integration of perception and action. This method addresses the limitations of previous models by enhancing functional competence and facilitating real-time interaction and adaptation.
@@
Language-driven object detection. Traditionally, language-driven object detection methods were designed for limited categories, a.k.a. closed-vocabulary object detection, like DETR [15]. Recently, using large vision-language models (VLMs) like CLIP [16] has opened doors to open-vocabulary object detection. Unlike traditional methods, open-vocabulary object detection methods OWLv2 [5] and RegionCLIP [6] are not restricted to predefined categories. The open-vocabulary object detection allows them to excel in tasks like zero-shot object detection and segmentation based on natural language descriptions. This shift towards open-vocabulary object detection [5] extends beyond just identifying objects. It has applications in related areas like affordance detection [17], where the focus is identifying potential actions that can be performed on objects by the robot. Inspired by OWLv2, we utilize this method in our framework to enable zero-shot object detection via natural language commands.
@@
Object pose localization. Recent advancements in object pose estimation have focused on developing techniques that handle novel objects without requiring retraining, addressing a key challenge in applying deep learning models to industrial settings. Nguyen et al. [18] introduced GigaPose, simplifying the pose estimation process to a single correspondence match between a query image and a set of templates. Recently, Wen et al. [7] presented FoundationPose, a unified 6D object pose estimation and tracking model, utilizing large-scale synthetic training data and a novel transformer-based architecture. FoundationPose supports both model-based and model-free setups with novel objects. However, manual segmentation of objects in the first frame from pre-recorded videos is still required. In this work, we integrate a language-driven object detection module that could enable end-to-end applications from human commands to robotic grasping actions in a closed-loop system.
@@
Real-time trajectory optimization. In recent years, significant interest has been in extending trajectory optimization to online planning using a receding horizon. This interest spans both gradient-based methods [19] and sampling-based approaches [20]. However, point-to-point planning is often not enough for manipulation tasks. Additional constraints, such as pre-grasp points, must be considered too. The sequence-of-constraints Model Predictive Control (MPC) introduced by Toussaint et al. [21] addresses task-and-motion planning (TAMP) in three phases: initially, task planning generates waypoints; subsequently, the timing of these waypoints is optimized to form a reference trajectory; finally, MPC tracks this reference trajectory to compute collision-free paths over a short planning horizon. Note that this approach requires a global reference to incorporate waypoints into the MPC. Tackling this issue, in our recent work [8], we propose the model-predictive trajectory optimization (MP-TrajOpt) that avoids calculating a reference trajectory for waypoint timing. Instead, this MPC approach employs a cost-to-go analysis for waypoints. The advantages of integrating MP-TrajOpt are twofold. First, to complete a general grasping task, the robot must navigate specific waypoints to avoid collisions between its gripper and the object. For example, in a top-down grasp, the robot should pass through a waypoint located above the object along the -axis, as illustrated in Figure 3. Second, passing waypoints with non-zero velocity allows for a more natural and smooth grasping action, which differs from most state-of-the-art methods [22, 23].
@@
Leveraging the vast knowledge from foundation models, we tackle the language-driven tasks in this world. Our contributions are summarized as follows:
@@ III Methods
This section explains the three main modules: language-driven object detection, real-time object pose localization, and online trajectory optimization (MP-TrajOpt), illustrated in Figure 1. Consider a scenario in a workshop where a human instructs a robot to grasp a tool with a known CAD model and place it in a predefined location. First, the language-driven object detection module processes the image to determine the 2D location of the tool and generate a binary mask. This mask, combined with the CAD model, is used by the real-time object pose localization module to estimate the initial pose of the tool. A Kalman filter refines the object’s pose for smoothness. Finally, the online trajectory optimization module plans and executes the grasp and placement actions, ensuring smooth and collision-free movements.
@@ III-A Language-Driven Object Detection
To determine the 2D location of a target object specified in a natural language command, such as ”Grasp the orange drill,” we first extract key phrases from the command, in this case, ”orange drill.” Concurrently, an RGB image is input into the OWLv2 model [5] to generate pseudo-box annotations for image-text pairs, as illustrated in Figure 2. Although we utilize the pre-trained weights of the OWLv2 model without additional training, this approach is efficient and effective at pinpointing the object’s center within the image. In comparison to more complex segmentation methods, such as SAM [24] or CNOS [25], which require several seconds for inference, the OWLv2 model can ground the object within across all experimental cases. A binary mask is generated upon obtaining the target object’s bounding box. This binary mask serves as the input for the subsequent object pose localization process.
@@ III-B Object pose localization module
In the pose initialization stage, we align and synchronize data from the depth sensor and the RGB sensor of an RGB-D camera. Subsequently, leveraging the object detection mask generated by the method mentioned earlier, we formulate an initial coarse pose estimation. This initial estimation is the input to a pose refinement network in the subsequent stage. The architecture of this refinement network, depicted in the second stage of Figure 2, is designed to infer translation updates and rotation updates of the corresponding object. Each update is individually processed through a transformer encoder[26] and then linearly projected to the output dimension. Both and are expressed relative to the camera frame. The initial coarse pose is refined utilizing the following update equations [7]:
where indicates the update operation on . This disentangled representation approach effectively decouples the translation update from the orientation update. The refinement network is trained under supervision [7] using the L2 loss function:
@@
To enhance the accuracy and smoothness of pose estimation ] of the object, we incorporate a Kalman Filter (KF) [27] of the state vector
with is the vector of three Euler angles computed from the object’s orientation [28]. We consider the process model for the KF with the update time and at the time index in the following
@@ III-C1 Mathematical model
with the generalized coordinates , the positive definite mass matrix , the matrix of Coriolis and centrifugal terms , the gravitational forces , and the generalized torques .
@@ III-C3 Planning
where the cost terms and drive the trajectory to the waypoint and the goal, respectively. The regularization cost on ensures smooth trajectories, and is a collision avoidance term. The weights are tuning parameters. The trajectory must adhere to the system dynamics (15b) with the initial states obtained by the previous planner iteration (15c). The endpoint is constraint to an equilibrium point in (15d), and the states and inputs are bounded using upper , and lower , bounds. The endpoints of the split horizon are constrained to a waypoint tolerance set and a goal tolerance set , respectively. For details on computing the horizon lengths and , refer to [8].
We use a state machine to guide the robot through the pick-and-place task. First, the robot must approach the object to grasp it. We derive a grasp pose and a pre-grasp waypoint from the object pose. The pre-grasp pose is placed above the object to ensure a proper approach direction and a collision-free grasp. Figure 3 illustrates an example of this pre-grasp waypoint. Note that moving the object further than a tolerance dynamically changes these poses. When the robot end-effector reaches the grasp pose, the robot grasps the object and can plan toward the place location. Similarly to the grasp poses, a pre-place and place pose are defined to place the object.
@@ IV Experiments
We experimentally evaluate the proposed framework on grasping tasks in real-world scenarios with a 7-DoF KUKA LBR iiwa 14 R820 and novel objects, captured in Figure 4. We utilize the D435i RealSense camera in an eye-to-hand setup where this camera observes the robot within its workspace like in Figure 3.
Three modules of the proposed framework, see Figure 1, are executed on a desktop PC with an Intel Core i7 10700K CPU and 32 GB of RAM. Inference processes of the language-driven module and the object pose localization run on the NVIDIA RTX 3080 GPU with 8GB of VRAM. The optimization problem described in Equation (15) is implemented as a Python-based ROS node utilizing CasADi [31]. It is solved using the nonlinear interior point solver IPOPT [32] and the MA57 linear solver, all running on an Ubuntu 22.04 system. This implementation achieves planning times under , which includes the online resolution of the analytic inverse kinematics for new Cartesian waypoints and a desired goal [30]. The solution nearest to the previous waypoint is selected in the inverse kinematics using a least-squares criterion for consistency.